What average acceleration does it take to stop a car traveling at 60 kph in 8 seconds
Answer by richwmiller(17219)

v=vinit+acceleration *time
0=60+acc*8
-60=acc*8
acc=-60/8
acc=-30/4
acc=-15/2
acc=-7.5 kps
which of course is deceleration